I have a long word problem that I need help with.
In 2000, crude oil production by OPEC countries was 29.2 million barrels per day. In 2002, OPEC crude oil production had risen to about 28.7 million barrels per day.
a)write two ordered pairs of the form (years after 2000, crude oil production) for this situation.
b)assume that the relationship between years after 1997 and crude oil production is linear over this period. Use the ordered pairs from part (a) to write an equation of the lines relating year to crude oil production.
c)Use the linear equation from part (b) to estimate the crude oil production by OPEC countries in 2001.
I am not sure how to work this problem. or where to start.
thanks
megan
: I have a long word problem that I need help with.
In 2000, crude oil production by OPEC countries was 29.2 million barrels per day. In 2002, OPEC crude oil production had risen to about 28.7 million barrels per day.
a)write two ordered pairs of the form (years after 2000, crude oil production) for this situation.
b)assume that the relationship between years after 1997 and crude oil production is linear over this period. Use the ordered pairs from part (a) to write an equation of the lines relating year to crude oil production.
c)Use the linear equation from part (b) to estimate the crude oil production by OPEC countries in 2001.
I am not sure how to work this problem. or where to start.
thanks
megan

Answer by ankor@dixie-net.com(4531)

n 2000, crude oil production by OPEC countries was 29.2 million barrels per day. In 2002, OPEC crude oil production had risen to about 28.7 million barrels per day.
:
OK, common sense tells us the oil production to has rise
lets reverse the values and proceed from there.
The method should be helpful at least.
:
In 2000, crude oil production by OPEC countries was 28.7 million barrels per day. In 2002, OPEC crude oil production had risen to about 29.2 million barrels per day.
;
a)write two ordered pairs of the form (years after 2000, crude oil production) for this situation.
:
Let x = years after 2000
Let y = production of crude in millions of barrels
:
Assign the given values as follows:
x1 = 0; y1 = 28.7
x2 = 2; y2 = 29.2
Find the slope (m) using the slope formula: m = (y2-y1)/(x2-x1)
m = (29.2 - 28.7)/(2 - 0) = .5/2 = +.25 is the slope
:
:
b)assume that the relationship between years after 1997 and crude oil production is linear over this period. Use the ordered pairs from part (a) to write an equation of the lines relating year to crude oil production.
:
Write the equation using the point/slope formula: y - y1 = m(x - x1)
Using our coordinates we have:
y - 28.7 = .25(x - 0)
y - 28.7 = .25x
y = .25x + 28.7 is the equation
:
1997 is 3 yrs before 2000, therefore x = -3 for 1997
:
Substitute -3 for x in our equation:
y = .25(-3) + 28.7
y = -.75 + 28.7
y = 27.95 million barrels in 1997
:
c)Use the linear equation from part (b) to estimate the crude oil production by OPEC countries in 2001.
2001 is one year after 2000, therefore x = 1
y = .25(1) + 28.7
y = 28.95 million barrels in 2001
:
:
A graph of this would be where x = 0 for the year 2000
 graph( 300, 200, -10, 10, -10, 35, .25x+28.7)